OpenStudy (anonymous):

. In triangle ABC, AB = 14, AC = 22 and angle A measures 43 degrees. Calculate the area of triangle ABC, correct to the nearest whole unit, and find the length of side BC, correct to the nearest tenth.

OpenStudy (anonymous):

i need help with this question plese

OpenStudy (anonymous):

area is \[\frac{1}{2}14\times 22\times \sin(43)\]

OpenStudy (anonymous):

thanks but can u plese show work thanks

OpenStudy (anonymous):

the work is from the formula that says area of a triangle is one half base times height.

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks sooo much the drawing makes it soo much esaier to understand

OpenStudy (anonymous):

since \[\sin(43)=\frac{h}{14}\] we know \[h=14\times \sin(43)\] and so \[\frac{1}{2}bh =\frac{1}{2}22\times 14\times \sin(43)\]

OpenStudy (anonymous):

for the second part of the question you need the law of cosines

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

yup thanks sooo much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw